A body with mass density 0.8g/cm cube and volume 100 cm cube is on the fluid with density 1.2 g/cm cube . Find mass of the body and amount of the volume inside the fluid.
1
Expert's answer
2020-05-25T11:18:17-0400

By definition of the density, the mass of the body will be:


"m = \rho V"

where "\rho" is the density and "V" is the volume of the body. Thus:


"m = 0.8 g/cm^3\cdot100cm^3 = 80g"

In equlibrium position forces that are acting on the body are equal. Thus:


"F_g = F_a"


where "F_g = mg" is the force of gravity and "F_a" is the Archimedes' force.


"mg = \rho_{lic}gV"

where "\rho_{lic} =1.2g/cm^3" is the density of the fluid and "V" is the volume of the body inside the fluid. Expressing "V" :


"V = \dfrac{mg}{\rho_{lic}g} = \dfrac{m}{\rho_{lic}} = \dfrac{80g}{1.2g/cm^3} = 66.7 cm^3"

Answer. m = 80 g, V = 66.7 cm^3.